4. Find the square roots of the following numbers by the Prime Factorisation Method.  vii. 5929 viii. 9216 ix. 529 x. 8100 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



5929 = 7×7×11×11
⇒ 5929 =(7×7)×(11×11)
⇒ 5929 =(7×11)×(7×11)
⇒ 5929 = (7×11)2
⇒√5929=7×11=77


9216 = 2×2×2×2×2×2×2×2×2×2×3×3×1
⇒9216=(2×2)×(2×2)×(2×2)×(2×2)×(2×2)×(3×3)
⇒9216=(2×2×2×2×2×3)×(2×2×2×2×2×3)
⇒9216=96×96
⇒√9216=96


529 = 23×23 529
√529 = 23


8100 = 2×2×3×3×3×3×5×5×1
⇒ 8100 = (2×2) ×(3×3)×(3×3)×(5×5) ⇒ 8100 = (2×3×3×5)×(2×3×3×5)
⇒ 8100 = 90×90
⇒ 8100 = (90)2
⇒√8100=90


Thanks